The shaft of an elevator motor turns clockwise at 180 rpm for 10 s, then turns counter-clockwise at 240 rpm for 5s. What is the angular displacement of the shaft during this motion? answer in radians please!
Answer #1

for the first 10 seconds

angular velocity = 180rpm = 180 x 2π / 60 = 18.85 rad/s

so angular displacement = 18.85 x 10 = 188.5 radian (clockwise)

for later 5 seconds

angular velocity = 240rpm = 240 x 2π / 60 = 25.13 rad/s

so angular displacement = 25.13 x 5 = 125.55 radian (counterclockwise)

so net angular displacement = 188.5 - 125..55 = 62.83 radians clockwise